OpenStudy (anonymous):

twice a number plus four equals 6. what is the number?

OpenStudy (anonymous):

(6-4)/2

OpenStudy (anonymous):

2n+4=6 n=1

jimthompson5910 (jim_thompson5910):

"twice a number plus four equals 6" ---> 2x+4=6 2x+4=6 2x=6-4 2x=2 x=2/2 x=1 So the number is 1

OpenStudy (anonymous):

thank you guys!